# Data-Driven Reachability with Scenario Optimization and the Holdout Method

### v1

@@ IV-B Reachable Tubes
We define a forward estimated reachable tube as ^$\hat{\mathcal{R}}(t)=\{\Phi(t_{i};t_{0},x_{0},d):\forall t_{i}\in[t_{0},t],x_{0}\in X_{0},d\in D\}$ where $[t_{0},t]$ is a finite time interval, $X_{0}\subseteq\mathbb{R}^{n_{x}}$ is the set of initial states, $D$ is the set of disturbance signals $d:[t_{0},t]\rightarrow\mathbb{R}^{n_{d}}$, and $\Phi:\mathbb{R}\times X_{0}\times D\rightarrow\mathbb{R}^{n_{x}}$ is the state transition function. ^$\hat{\mathcal{R}}(t)$ is the set of all states to which the system can transition within a duration of time $[t_{0},t]$ for finite time steps $t_{i}$ from $X_{0}$ subject to disturbances in $D$. The reachable tube accounts for all time steps in interval $[t_{0},t]$, while the reachable sets presented in Section II only consider one specific time instant. We generalize the nonconvex scenario reachability method presented above and pose the problem of finding a reachable tube directly from data as a chance-constrained optimization problem with regularization over time. In particular, we propose an approach in which we construct a smoothed, finite set of time-varying RBFs:
@@
The set of initial states is the interval such that $x_{1}(0),x_{2}(0)\in[1,1.25]$, and we take $\mu_{X_{0}}$ to be the uniform random variable over this interval. The time range is $[t_{0},t]=[0,10]$. We calculate the reachable tube estimate using one radial basis function for every time instant. To apply the holdout method, we partition our data set of trajectories into a training and test set, as was done in the previous section, and take $\beta=10^{-9}$. We consider a boundary violation to be any trajectory that falls outside the reachable tube at any time instant.
@@ V De-randomization
In this work, we have advocated for the use of a general purpose, practical, and sharp method for data-driven reachability: the holdout method. The probabilistic guarantee offered by data-driven reachability in general comprises two “layers” of probability (cf. Theorem 1): the bound on the violation probability itself, and the bound in probability, with respect to the probability law of the sample generation procedure. In other words, the confidence that the samples have yielded enough information to construct a reachable set estimate for which the violation probability bound holds. These types of Probably Approximately Correct (PAC) bounds [32] are ubiquitous in statistical learning theory and data-driven approaches to probabilistic verification. They are generally recognized as the strongest type of guarantee that can be made while placing minimal assumptions on the random variable being verified. Nevertheless, it is reasonable to object to two layers of probability, both on the grounds of interpretability and soundness, compared to conventional control-theoretic guarantees, which give a guarantee of 100% certainty, subject to accurate modeling information. To that end, several works have made attempts to “de-randomize” data-driven methods, augmenting them with side information to remove one or both of the probabilistic layers [33, 34, 35].
In this section we argue that, while de-randomization aligns with control theory’s pursuit of certainty, it is not practical for the type of data-driven method explored in this paper and similar approaches. Data-driven verification has two primary advantages over other methods: mild requirements on available side information, often times requiring none at all, and computational efficiency compared to deterministic sampling-based methods. Given the considerable sacrifice in the certitude that one must make in admitting a probabilistic safety guarantee, the loss of either advantage is inadmissible. We have found that de-randomized data-driven methods often, and perhaps by necessity, relinquish both.
@@ V-A De-randomization Methods
The removal of either layer of probability (or both) is carried out by selecting, according to some side information, a suitable enlargement of the data-driven estimator. For simplicity of exposition, we will focus on de-randomizing the inner probability layer, and return to the outer layer in a moment. Specifically, suppose we have found a feasible point $\theta^{*}$ satisfying the chance constraint $\mathbf{P}\{g(\theta^{*},x)\leq 0\}\geq 1-\epsilon$. To de-randomize such a bound is to establish some $\gamma\geq 0$, as a function of information not available to the original data-driven algorithm, such that $g(\theta^{*},x)\leq\gamma$ holds almost surely.
A general scheme for how to perform a de-randomization of this type is laid out in [34]. The crux of the scheme is computing a uniform level-set bound function $h(\epsilon)$ that acts as a uniform upper bound (over $\theta$) on the quantile function of the random variable $(\sup_{x}g(x,\theta)-g(x,\theta))$; as a consequence, it follows that if $\mathbf{P}\{g(\theta^{*},x)\leq 0\}\geq 1-\epsilon$, then $\mathbf{P}\{g(\theta^{*},x)\leq h(\epsilon)\}=1$. In our case, the function $g$ contains information about the dynamical system specified in the reachability problem, so de-randomization requires some amount of new side information.
A standard assumption is to presume knowledge of a global Lipschitz bound on the dynamics and to take the estimator sets to be $p$-norm balls, or ellipsoids for simplicity. In this case, the magnitude of the enlargement factor from existing work scales exponentially with state dimension [34, cf. Remark 3.9], [33, cf. Theorem 1]. Interestingly, a less conservative estimate of the reachable set with the same number of queries—and moreover with a non-stochastic guarantee—is possible with a “sample and cover” approach that spreads the query points uniformly over the initial set and uses a contraction-based approximation to completely cover a region around each query point. Given this context, it is natural to wonder if the situation can be improved with a sharper analysis, or if there is a fundamental limitation with this type of de-randomization in general.
@@ V-B Lower Bounds from Zeroth-Order Optimization
Rather than being an artifact of the analysis, the exponential scaling mentioned previously is inherent to any approach, to ensure zero violation probability using only sampling and a Lipschitz condition. To demonstrate this, suppose we are given an $h:B_{2}^{d}\mapsto\mathbb{R}$, where $B_{2}^{d}(r):=\{x\in\mathbb{R}^{d}\mid\|x\|_{2}\leq r\}$ and $B_{2}^{d}:=B_{2}^{d}(1)$, and we are assured that (a) $h$ is $L$-Lipschitz and (b) the following probabilistic guarantee holds:
@@
The implication of Lemma 1 is as follows. Supposing that the $\varepsilon$ from Lemma 1 decreases at a rate of $1/M$, where $M$ is the number of holdout examples (cf. the discussion after Theorem 1 in Section III), one needs at least $(L/\gamma)^{d}$ samples to ensure a bound $\max_{x\in B_{2}^{d}}h(x)\leq\gamma$ in this setting.
@@
In contrast, data-driven methods that yield PAC bounds require almost no side information and are quite sample efficient, as demonstrated in this work. Since the costs incurred by de-randomization are large relative to this baseline, we argue that whenever a sampling-based approach is deemed acceptable or necessary, the circumstances where de-randomization is worth the price are rare.
@@ VII Acknowledgments
This paper is supported in part by the NSF project CNS-2111688. The first author was also supported by a NSF Graduate Research Fellowship.

### v2

@@ IV-B Reachable Tubes
We define a forward estimated reachable tube as $\hat{\mathcal{R}}(t)=\{\Phi(t_{i};t_{0},x_{0},d):\forall t_{i}\in[t_{0},t],x_{0}\in X_{0},d\in D\}$ where $t_{0},t]$ is a finite time interval, $X_{0}\subseteq\mathbb{R}^{n_{x}}$ is the set of initial states, $D$ is the set of disturbance signals $d:[t_{0},t]\rightarrow\mathbb{R}^{n_{d}}$, and $\Phi:\mathbb{R}\times X_{0}\times D\rightarrow\mathbb{R}^{n_{x}}$ is the state transition function. $\hat{\mathcal{R}}(t)$ is the set of all states to which the system can transition within a duration of time $t_{0},t]$ for finite time steps $t_{i}$ from $X_{0}$ subject to disturbances in $D$. The reachable tube accounts for all time steps in interval $t_{0},t]$, while the reachable sets presented in Section II only consider one specific time instant. We generalize the nonconvex scenario reachability method presented above and pose the problem of finding a reachable tube directly from data as a chance-constrained optimization problem with regularization over time. In particular, we propose an approach in which we construct a smoothed, finite set of time-varying RBFs:
@@
The set of initial states is the interval such that $x_{1}(0),x_{2}(0)\in[1,1.25]$, and we take $\mu_{X_{0}}$ to be the uniform random variable over this interval. The time range is $t_{0},t]=[0,10]$. We calculate the reachable tube estimate using one radial basis function for every time instant. To apply the holdout method, we partition our data set of trajectories into a training and test set, as was done in the previous section, and take $\beta=10^{-9}$. We consider a boundary violation to be any trajectory that falls outside the reachable tube at any time instant.
@@ V De-randomization
In this work, we have advocated for the use of a general purpose, practical, and sharp method for data-driven reachability: the holdout method. Data-driven reachability offers probabilistic guarantees that are comprised of two “layers” of probability (cf. Theorem 1): the bound on the violation probability, and the bound in probability, with respect to the probability law of the sample generation (in other words, the confidence that the samples have yielded enough information to construct a reachable set estimate for which the violation probability holds). These types of Probably Approximately Correct (PAC) bounds [32] are ubiquitous in statistical learning theory and data-driven approaches to probabilistic verification. They are generally recognized as the strongest type of guarantee that can be made while placing minimal assumptions on the random variable verified. Nevertheless, it is reasonable to critize two layers of probability on the grounds of interpretability and soundness, compared to conventional control-theoretic guarantees, which provide 100% certainty given accurate modeling information. To that end, several works have made attempts to “de-randomize” data-driven methods, augmenting them with additional information to remove one, or both, of the probabilistic layers [33, 34, 35].
In this section we argue that, while de-randomization aligns with control theory’s pursuit of certainty, it is not practical for the type of data-driven method explored in this paper and similar approaches. Data-driven verification has two primary advantages: mild requirements on side information, often requiring none at all, and computational efficiency compared to deterministic, sampling-based methods. We have found that de-randomized data-driven methods often relinquish both advantages.
@@ V-A De-randomization Methods
The removal of either layer of probability (or both) is carried out by selecting, according to some side information, a suitable enlargement of the data-driven estimator. For simplicity, we first focus on de-randomizing the inner probability layer, and return to the outer layer later. Specifically, suppose we have found a feasible point $\theta^{*}$ satisfying the chance constraint $\mathbf{P}\{g(\theta^{*},x)\leq 0\}\geq 1-\epsilon$. To de-randomize such a bound is to establish some $\gamma\geq 0$, as a function of information not available to the original data-driven algorithm, such that $g(\theta^{*},x)\leq\gamma$ holds almost surely.
A general scheme for how to perform a de-randomization of this type is laid out in [34]. The central idea of the scheme is computing a uniform level-set bound function $h(\epsilon)$ that acts as a uniform upper bound (over $\theta$) on the quantile function of the random variable $(\sup_{x}g(x,\theta)-g(x,\theta))$. As a consequence, it follows that if $\mathbf{P}\{g(\theta^{*},x)\leq 0\}\geq 1-\epsilon$, then $\mathbf{P}\{g(\theta^{*},x)\leq h(\epsilon)\}=1$. In our case, the function $g$ contains information about the dynamical system specified in the reachability problem, so de-randomization requires some amount of new side information.
A standard assumption is to presume knowledge of a global Lipschitz bound on the dynamics and to take the estimator sets to be $p$-norm balls, or ellipsoids. In this case, the magnitude of the enlargement factor from existing work scales exponentially with state dimension [34, cf. Remark 3.9], [33, cf. Theorem 1]. Interestingly, a less conservative estimate of the reachable set with the same number of queries—and moreover with a non-stochastic guarantee—is possible with a “sample and cover” approach. This approach spreads query points uniformly over the initial set and uses a contraction-based approximation to completely cover a region around each query point. Given this context, it is natural to wonder if the situation can be improved with a sharper analysis, or if there is a fundamental limitation with this type of de-randomization in general.
@@ V-B Lower Bounds from Zeroth-Order Optimization
The exponential scaling mentioned previously is inherent to all de-randomization approaches — ensuring zero violation probability using only samples and a Lipschitz condition. To demonstrate this, suppose we are given an $h:B_{2}^{d}\mapsto\mathbb{R}$, where $B_{2}^{d}(r):=\{x\in\mathbb{R}^{d}\mid\|x\|_{2}\leq r\}$ and $B_{2}^{d}:=B_{2}^{d}(1)$, and we are assured that (a) $h$ is $L$-Lipschitz and (b) the following probabilistic guarantee holds:
@@
The implication of Lemma 1 is as follows. Suppose that the $\varepsilon$ from Lemma 1 decreases at a rate of $1/M$, where $M$ is the number of holdout examples, one needs at least $(L/\gamma)^{d}$ samples to ensure a bound $\max_{x\in B_{2}^{d}}h(x)\leq\gamma$ in this setting.
@@
In contrast, data-driven methods that yield PAC bounds require almost no side information and are quite sample efficient, as demonstrated in this work. Since the costs incurred by de-randomization are large relative to this baseline, we argue that whenever a sampling-based approach is deemed acceptable, the circumstances where de-randomization is worth the price are rare.
@@ VII Acknowledgments
This paper is supported in part by the NSF project CNS-2111688. The first author was also supported by an NSF Graduate Research Fellowship.